OpenStudy (donnie1999):

------------WILL FAN AND MEDAL------------ You are serving bratwurst and hamburgers at your annual picnic. You want at least three bratwursts or hamburgers for each of your 20 guests. Bratwursts cost $1.2 each and hamburgers $0.95 each. Your budget is $75. Let x be the number of bratwursts and y the number of hamburgers. Which system of inequalities represents this situation?

OpenStudy (donnie1999):

@pinkbubbles

OpenStudy (anonymous):

I will do this

OpenStudy (anonymous):

1.2 as $1.20?

OpenStudy (anonymous):

What are the answer choices?

OpenStudy (anonymous):

It would come to 43

OpenStudy (anonymous):

$

OpenStudy (anonymous):

Did you get it right?

OpenStudy (anonymous):

@donnie1999

OpenStudy (donnie1999):

x+y>=60 1.2x+0.95y<=75 x+y<=20 1.2x+0.95y>=75 x+y<=20 1.2x+0.95y<=75 x+y>=60 1.2x+0.95y>=75

OpenStudy (anonymous):

x+y=>60 1.2x+0.95y<=75